# Automorphisms of relatively hyperbolic groups and the Farrell–Jones Conjecture

### 23 Nov 2023 v1

@@ 5.B. The Periodic JSJ tree
By Lemma 5.14, up to taking a power of $\Phi_{H}$, we may suppose that $F_{\Phi_{H}}$ fixes $v$. Let $p_{\phi}(v)$ be the projection of $v$ in $T^{\phi}$. We distinguish between several cases, according to the nature of $p_{\phi}(v)$. Note that, by Corollary 5.9, the vertex $p_{\phi}(v)$ is not QH with fibre.
Suppose that $p_{\phi}(v)$ is elementary. Then $\operatorname{\mathrm{Stab}}(v)$ is elementary. Thus, $\operatorname{\mathrm{Stab}}(v)$ splits naturally as $\operatorname{\mathrm{Stab}}(v)=\operatorname{\mathrm{Stab}}(v)\ast_{H\cap\operatorname{\mathrm{Stab}}(v)}(H\cap\operatorname{\mathrm{Stab}}(v))$ and $H\cap\operatorname{\mathrm{Stab}}(v)$ is elementary.
Suppose now that $p_{\phi}(v)$ is rigid. By Lemma 5.4 applied to $p_{\phi}(v)$, up to taking a power of $\Phi_{H}$, we may suppose that $\Phi_{H}$ acts on $\operatorname{\mathrm{Stab}}(p_{\phi}(v))$ as a global conjugation by an element $g\in\operatorname{\mathrm{Stab}}(p_{\phi}(v))$.
If $g$ is finite, up to taking a power of $\Phi_{H}$, we see that $\Phi_{H}$ acts as the identity on $\operatorname{\mathrm{Stab}}(p_{0}(v))$, so that $\operatorname{\mathrm{Stab}}(v)\subseteq\operatorname{\mathrm{Stab}}(p_{0}(v))\subseteq H$.
So we may suppose that $g$ is infinite. Then $H\cap\operatorname{\mathrm{Stab}}(p_{\phi}(v))$ consists of all elements $h\in\operatorname{\mathrm{Stab}}(p_{\phi}(v))$ which commute with a power of $g$. But the commensurator of $\langle g\rangle$ in $G$ is elementary (because maximal elementary subgroups are almost malnormal, see [GL15, Corollary 3.2]). Thus, the group $H\cap\operatorname{\mathrm{Stab}}(p_{\phi}(v))$ is elementary. Hence the group $H\cap\operatorname{\mathrm{Stab}}(v)$ is elementary and the splitting $\operatorname{\mathrm{Stab}}(v)=\operatorname{\mathrm{Stab}}(v)\ast_{H_{j}\cap\operatorname{\mathrm{Stab}}(v)}(H_{j}\cap\operatorname{\mathrm{Stab}}(v))$ is over an elementary subgroup.
As we have ruled out every case, this concludes the proof. ∎
Lemma 5.18.
Suppose that $\phi$ is almost rotationless. For every $H\in\mathrm{Per}_{\mathrm{NP}}(\phi)$, and the unique vertex $v_{H}\in T^{\mathrm{Per}}$ it stabilises, we have $\operatorname{\mathrm{Stab}}(v_{H})=H$.
@@
Let $H\in\mathrm{Per}_{\mathrm{NP}}(\phi)$. By Lemma 5.16, it suffices to show that the stabiliser $G_{T^{\mathrm{ref}}_{H}}$ of $T^{\mathrm{ref}}_{H}$ is equal to $H$.
Claim 1. The stabiliser of any edge of $T^{\mathrm{ref}}_{H}$ is contained in $H$.
Proof..
Let $e\in ET^{\mathrm{ref}}_{H}$. Note that $e$ is contained in the axis of an element $g\in H$. Since $H$ is elliptic in $T^{\mathrm{Per}}$, the image $p_{\operatorname{Per}}(e)$ of $e$ in $T^{\mathrm{Per}}$ is collapsed to a point.
By minimality of $T^{\mathrm{ref}}$, the image $p_{\phi}(e)$ is an edge, and $G_{e}=G_{p_{\phi}(e)}$. By Corollary 5.8, any edge contained in $T^{\phi}_{H}$ is adjacent to a rigid vertex $v$, and $G_{p_{\phi}(e)}\leq G_{v}\leq H$. ∎
Claim 2. Let $v\in VT^{\mathrm{ref}}_{H}$. The stabiliser $\operatorname{\mathrm{Stab}}(v)$ of $v$ in $G$ is contained in $H$.
Proof..
Suppose towards a contradiction that $\operatorname{\mathrm{Stab}}(v)$ is not contained in $H$. By Lemma 5.17, there exists of splitting $S_{v}$ of $\operatorname{\mathrm{Stab}}(v)$ with elementary edge stabilisers such that $H\cap\operatorname{\mathrm{Stab}}(v)$ fixes a point $w$ in $S_{v}$ distinct from any point fixed by $\operatorname{\mathrm{Stab}}(v)$. This splitting induces a refinement $T^{\prime}$ of $T^{\mathrm{ref}}$ by blowing up $S_{v}$ at the vertex $v$ and attaching the adjacent edges accordingly. We make the additional assumption that if $e$ is an edge in $T^{\mathrm{ref}}$ adjacent to $v$ such that $\operatorname{\mathrm{Stab}}(e)\subseteq\operatorname{\mathrm{Stab}}(w)$, then we attach $e$ to $w$. The tree $T^{\prime}$ is an $({\mathcal{A}},{\mathcal{P}})$-tree since any edge of $T^{\mathrm{ref}}$ or $S_{v}$ is elementary and since $\operatorname{\mathrm{Stab}}(v)$ fixes a point in $S_{v}$ if $v$ is elementary by Lemma 5.17.
By Claim $1$, every edge of $T^{\mathrm{ref}}$ is attached to $w$ when blowing up the splitting $S_{v}$. Therefore, the minimal tree $T_{H}^{\prime}$ of $H$ in $T^{\prime}$ does not contain any vertex stabilised by $\operatorname{\mathrm{Stab}}(v)$. Collasping, for every $K\in\mathrm{Per}_{\mathrm{NP}}(\phi)$, the tree $T_{K}^{\prime}$ gives an $({\mathcal{A}},{\mathcal{P}}\cup\mathrm{Per}_{\mathrm{NP}}(\phi))$-tree $U$ such that $\operatorname{\mathrm{Stab}}(v_{H})$ does not fix a point in it. This contradicts the fact that $\operatorname{\mathrm{Stab}}(v_{H})$ is elliptic in every $({\mathcal{A}},{\mathcal{P}}\cup\mathrm{Per}_{\mathrm{NP}}(\phi))$-tree (see Theorem 4.2 $(7)$). Hence, we have $\operatorname{\mathrm{Stab}}(v)\subseteq H$. ∎
Claim 3. Suppose that $g\in G_{T^{\mathrm{ref}}_{H}}$. Then $g\in H$.
Proof..
If $g$ is elliptic, then since $g$ preserves $T^{\mathrm{ref}}_{H}$, one of its fixed point is contained in $T^{\mathrm{ref}}_{H}$. By Claim $2$, we see that $g\in H$.
So we may suppose that $g$ is loxodromic in $T^{\mathrm{ref}}$. We claim that there exists $v\in VT^{\mathrm{ref}}_{H}$ with infinite nonelementary stabiliser. Indeed, first note that $T^{\mathrm{ref}}_{H}$ is not reduced to a point as $g$ is loxodromic. In particular, $H$ does not fix a point of $T^{\mathrm{ref}}$. Thus, by minimality of $T^{\mathrm{ref}}$, the projection $p_{\phi}(T^{\mathrm{ref}}_{H})$ is nontrivial. By Corollary 5.9, the tree $p_{\phi}(T^{\mathrm{ref}}_{H})$ contains a rigid vertex and the preimage in $T^{\mathrm{ref}}$ of such a rigid vertex contains a vertex $v$ with infinite nonelementary stabiliser.
Let $h\in\operatorname{\mathrm{Stab}}(v)$ be of infinite order. Since the action of $G$ on $T$ is acylindrical, up to taking a vertex $v$ far enough from the axis of $g$, we may suppose that $h$ and $ghg^{-1}$ do not have a common fixed point in $T$. In particular, they generate a nonabelian free group by standard ping pong arguments. Moreover, this nonabelian free group contains a finitely generated nonabelian free subgroup $L$ consisting of loxodromic elements, hence consisting of nonperipheral elements.
Note that $h,ghg^{-1},g^{2}hg^{-2}\in H$ since $g$ preserves $T^{\mathrm{ref}}_{H}$ and since any vertex stabiliser of $T^{\mathrm{ref}}_{H}$ is contained in $H$ by Claim $2$. Thus, up to taking a power of $\Phi$, we may suppose that $\langle L,gLg^{-1}\rangle\leq\langle h,ghg^{-1},g^{2}hg^{-2}\rangle\leq\operatorname{\mathrm{Fix}}(\Phi)$. Let ${\mathcal{K}}=\{[\langle L,gLg^{-1}\rangle]\}$ and let $T_{\mathcal{K}}$ be the canonical JSJ tree associated with ${\mathcal{K}}$ given by Theorem 4.2.
Since $\langle L,gLg^{-1}\rangle$ is nonelementary, it fixes a unique rigid point $w\in VT_{\mathcal{K}}$ by 4.2 $(9)$. We claim that $g$ fixes $w$. Indeed, otherwise the path between $w$ and $gw$ would be fixed by $gLg^{-1}$ which is a nonelementary subgroup. This would contradict Theorem 4.2 $(2)$.
Hence we have $g\in\operatorname{\mathrm{Stab}}(w)$. Since $w$ is rigid and since $\langle L,gLg^{-1}\rangle$ is finitely generated, by Theorem 4.2 $(10)$, the automorphism $\Phi$ acts as a global conjugation on $\operatorname{\mathrm{Stab}}(w)$. Since $\Phi$ fixes $L$ which is nonelementary, it must act as a periodic automorphism of $\operatorname{\mathrm{Stab}}(w)$. In particular, we have $g\in H$, which concludes the proof of the claim. ∎
Combining Lemma 5.16 and by Claim 3, we see that the stabiliser of $v_{H}$ is equal to $H$, which concludes the proof. ∎
Theorem 5.19.
Let $G$ be a one-ended hyperbolic group relative to ${\mathcal{P}}$ and let $\phi=[\Phi]\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}})$. Let $\phi^{N}$ be an almost rotationless power of $\phi$. For every rigid vertex $v\in VT^{\mathrm{Per}}$, there exists $[H]\in\mathrm{Per}_{\mathrm{NP}}(\phi^{N})$ such that $\operatorname{\mathrm{Stab}}(v)=H$. Conversely, for every $[H]\in\mathrm{Per}_{\mathrm{NP}}(\phi^{N})$, there exists a rigid vertex $v\in VT^{\mathrm{Per}}$ with $\operatorname{\mathrm{Stab}}(v)=H$.
Proof.
@@
Up to taking a power of $\phi$, we may assume that $\phi$ is almost rotationless (see Remark 5.13). Let $v\in VT^{\mathrm{Per}}$ be rigid. We claim that there exists $[H]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$ with $H\leq\operatorname{\mathrm{Stab}}(v)$. Consider the pre-image of $v$ in $T^{\mathrm{ref}}$. If it is a vertex, then $\operatorname{\mathrm{Stab}}(v)$ also stabilises a non-elementary vertex $w$ of $T^{\phi}$. If $w$ is rigid, then $\langle\phi\rangle^{0}\to\operatorname{\mathrm{Out}}(G_{v})$ has finite image, so $\operatorname{\mathrm{Stab}}(v)$ is periodic for some $\phi^{k}$, and so for $\phi$ (since we assume it is almost rotationless).
@@
Otherwise $w$ is a QH with fibre vertex. Note that all elementary subgroups of $\operatorname{\mathrm{Stab}}(w)$ are virtually cyclic. In particular, the edges adjacent to $v\in VT^{\mathrm{Per}}$ are all virtually cyclic, hence finitely generated. Thus, $\operatorname{\mathrm{Stab}}(v)$ is a finitely generated subgroup of $\operatorname{\mathrm{Stab}}(w)$. Since $w$ is a QH with fibre vertex, it is locally quasi-convex. Thus, $\operatorname{\mathrm{Stab}}(v)$ is hyperbolic.
@@
Note that, since $\operatorname{\mathrm{Stab}}(v)$ is hyperbolic and a subgroup of $\operatorname{\mathrm{Stab}}(w)$, the restriction ${\mathcal{P}}\cup\mathrm{Per}_{\mathrm{NP}}(\phi)|_{\operatorname{\mathrm{Stab}}(v)}$ of ${\mathcal{P}}\cup\mathrm{Per}_{\mathrm{NP}}(\phi)$ to $\operatorname{\mathrm{Stab}}(v)$ is a finite family of virtually cyclic subgroups of $\operatorname{\mathrm{Stab}}(v)$ by Lemma 5.2. Thus, we can apply [GL15, Theorem 3.9] to show that $\operatorname{\mathrm{Out}}(\operatorname{\mathrm{Stab}}(v),{\mathcal{P}}\cup\mathrm{Per}_{\mathrm{NP}}(\phi)|_{\operatorname{\mathrm{Stab}}(v)}\cup\mathrm{Inc}_{v})$ (which has $\operatorname{\mathrm{Out}}(\operatorname{\mathrm{Stab}}(v),{\mathcal{P}}\cup(\mathrm{Per}_{\mathrm{NP}}(\phi)|_{\operatorname{\mathrm{Stab}}(v)}\cup\mathrm{Inc}_{v})^{(t)})$ as a finite index subgroup) is infinite if and only if $\operatorname{\mathrm{Stab}}(v)$ has an $({\mathcal{A}},{\mathcal{P}}\cup\mathrm{Per}_{\mathrm{NP}}(\phi)|_{\operatorname{\mathrm{Stab}}(v)}\cup\mathrm{Inc}_{v})$-splitting. Since $v$ is rigid, no such splitting of $\operatorname{\mathrm{Stab}}(v)$ exists. Therefore, $\operatorname{\mathrm{Out}}(\operatorname{\mathrm{Stab}}(v),{\mathcal{P}}|_{\operatorname{\mathrm{Stab}}(v)}\cup\mathrm{Per}_{\mathrm{NP}}(\phi)|_{\operatorname{\mathrm{Stab}}(v)}\cup\mathrm{Inc}_{v})$ is finite. In particular, the group $\operatorname{\mathrm{Stab}}(v)$ is a nonelementary periodic subgroup of some power of $\phi$.
@@
Now suppose that the pre-image of $v$ contains an edge. Since $v$ is non-elementary, in fact this pre-image must be the minimal invariant tree $T^{\mathrm{ref}}_{\operatorname{\mathrm{Stab}}(v)}$. If this is contained in $T^{\mathrm{ref}}_{H}$ for some $H\in\mathrm{Per}_{\mathrm{NP}}(\phi)$, then $v$ is the unique vertex of $T^{\mathrm{Per}}$ stabilised by $H$, and by Lemma 5.18 $\operatorname{\mathrm{Stab}}(v)$ (as the stabiliser of this vertex) is equal to $H$.
@@
So suppose $T^{\mathrm{ref}}_{\operatorname{\mathrm{Stab}}(v)}$ is not contained in any $T^{\mathrm{ref}}_{H}$. Since the $T^{\mathrm{ref}}_{H}$ are disjoint by 5.15, there is some edge of $T^{\mathrm{ref}}_{\operatorname{\mathrm{Stab}}(v)}$ contained in no $T^{\mathrm{ref}}_{H}$, and hence collapsing all the $T^{\mathrm{ref}}_{H}$ will give an $({\mathcal{A}},{\mathcal{P}}\cup\mathrm{Per}_{\mathrm{NP}}(\phi))$ tree where $\operatorname{\mathrm{Stab}}(v)$ is not elliptic, contradicting 4.2(7).
@@
Conversely, let $[H]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$. By construction of $T^{\mathrm{Per}}$, the group $H$ fixes a vertex $v$ of $T^{\mathrm{Per}}$. Since $H$ is nonelementary, such a vertex is unique and is rigid by 4.2 $(10)$. By Lemma 5.18, we have $\operatorname{\mathrm{Stab}}(v)=H$. ∎
@@
Let $G$ be a one-ended hyperbolic group relative to ${\mathcal{P}}$ and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{P}})$. Let ${\mathcal{H}}=\bigcup_{n\in\mathbb{N}^{*}}\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$.
@@
Let $\gamma$ be a geodesic edge path of length $3$ in $T^{\mathrm{Per}}$ preserved by an automorphism $\Psi\in\phi$. Suppose that there exists an edge $e$ of $\gamma$ whose stabiliser is virtually cyclic (this applies in particular when one of the vertices of $\gamma$ is QH with fibre). Hence $\operatorname{\mathrm{Out}}(G_{e})$ is finite. Thus, $\Psi$ has a power acting as the identity on the infinite cyclic subgroup of $G_{e}$.
Thus, we may suppose that $\gamma$ only contains elementary and rigid vertices. Since $\gamma$ has length $3$ and since $T^{\mathrm{Per}}$ is bipartite, it contains an interior vertex $v$ which is rigid. Let $e_{1},e_{2}$ be the two edges of $\gamma$ adjacent to $v$. Then $G_{e_{1}}\cap G_{e_{2}}$ is finite by 4.2 $(8)$ and for every $i\in\{1,2\}$ the group $G_{e_{i}}$ is its own normaliser in $G_{v}$.
Let $\phi^{N}$ be an almost rotationless power of $\phi$. By 5.19, there exists $[H]\in\mathrm{Per}_{\mathrm{NP}}(\phi^{N})$ such that $\operatorname{\mathrm{Stab}}(v)=H$. Let $\Phi_{H}\in\phi^{N}$ be such that $\mathrm{Per}(\Phi_{H})=H$. Since $\Psi$ preserves $H$, and since $H$ is its own normaliser by Lemma 5.18, there exists $h\in H$ such that $\Psi^{N}=\mathrm{ad}_{h}\Phi_{H}$. As $\Psi^{N}$ preserves both $G_{e_{1}}$ and $G_{e_{2}}$, we see that $h\in N(G_{e_{1}})\cap N(G_{e_{2}})$, which is finite as explained above. Up to taking a power of $\Phi_{H}$ and $\Psi^{N}$, we may assume that both $\Phi_{H}$ and $\Psi^{N}$ act trivially on $N(G_{e_{1}})\cap N(G_{e_{2}})$. Taking further powers of $\Psi^{N}$ and $\Phi_{H}$ shows that there exists $M\in\mathbb{N}$ such that $\Psi^{M}=\Phi_{H}^{M}$.
Let $g\in H$ be infinite order, which exists since $H$ is nonelementary. Then $\Psi$ has a power which fixes $g$. This proves Assertion $(1)$.
@@
By Theorem 4.2 $(2)$, the action of $G$ on $T^{\mathrm{Per}}$ is acylindrical. Thus, by Lemma 4.3 $(3)$ (which we can apply by Assertion $(1)$), it suffices to prove that, for every $n\in\mathbb{N}^{*}$, every $\Psi\in\phi^{n}$ and every $g\in\operatorname{\mathrm{Fix}}(\Psi^{n})$, the element $g$ is elliptic in $T$.
Since $\phi$ is almost rotationless, in order to prove Assertion $(2)$, it suffices to prove that, for every $n\in\mathbb{N}^{*}$, every $\Psi\in\phi$ and every $g\in\operatorname{\mathrm{Fix}}(\Psi^{n})$, the element $g$ is elliptic in $T$.
@@
Suppose now that $g$ is nonperipheral. Suppose towards a contradiction that $g$ is loxodromic in $T^{\mathrm{Per}}$. Since $\Psi^{n}(g)=g$, the characteristic set of the isometry $F_{\Psi^{n}}$ contains the axis of $g$. By Lemma 4.3 $(2)$, up to taking a power of $\phi$ and changing the representative $\Psi$, the isometry $F_{\Psi^{n}}$ fixes pointwise the axis of $g$. By Lemma 4.3 $(1)$, up to taking a power of $\Psi$, the automorphism $\Psi$ fixes elementwise a nonabelian free group of loxodromic elements. In particular, since every peripheral element fixes a point in $T^{\mathrm{Per}}$, we see that $\Psi$ fixes a nonabelian free group of nonperipheral elements. Since $\phi$ is almost rotationless, for every $n\geq 1$, we have $\mathrm{Per}_{\mathrm{NP}}(\phi)=\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$, we see that $[\mathrm{Per}(\Psi)]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$. Thus, there exists $[H]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$ such that $g\in\operatorname{\mathrm{Fix}}(\Psi)\subseteq H$. In that case, the element $g$ is elliptic in $T^{\mathrm{Per}}$ by construction of $T^{\mathrm{Per}}$, a contradiction.

### 02 Mar 2026 v2

@@ 5.B. The Periodic JSJ tree
Let $H\in\mathrm{Per}_{\mathrm{NP}}(\phi)$, let $\Phi\in\phi$ be the automorphism satisfying $\operatorname{Per}(\Phi)=H$, and $F_{\Phi}$ the associated isometry of $T^{\mathrm{ref}}$. By Lemma 5.19, it suffices to show that the stabiliser $G_{T^{\mathrm{ref}}_{H}}$ of $T^{\mathrm{ref}}_{H}$ is equal to $H$. By ˜5.18 and the fact that $gT^{\mathrm{ref}}_{H}=T^{\mathrm{ref}}_{gHg^{-1}}$ for every $g\in G$, the group $G_{T^{\mathrm{ref}}_{H}}$ is equal to the normaliser $N(H)$ of $H$ in $G$.
Suppose first that $H$ fixes a (necessarily unique) vertex $v\in VT^{\mathrm{ref}}$. Then $v$ has nonelementary stabiliser. By Lemma˜5.6, the point $p_{\phi}(v)$ is the unique rigid vertex fixed by $H$ in $T^{\phi}$. Thus, $N(H)$ fixes $p_{\phi}(v)$. By Lemma˜5.6, we have $N(H)\subseteq G_{p_{\phi}(v)}\subseteq H$ and $H=N(H)$.
Suppose now that $H$ does not fix a vertex. By Lemma˜4.3, $H$ contains a loxodromic element, and so $T^{\mathrm{ref}}_{H}$ is an infinite subtree of $T^{\mathrm{ref}}$. Let $g\in N(H)$. Let $v_{1}$ and $v_{2}$ be vertices in $T^{\mathrm{ref}}_{H}$, at distance greater than the acylindricity constant $K$. Observe that the acylindricity assumption implies that there are only finitely many elements $g^{\prime}$ with the property that $gv_{i}=g^{\prime}v_{i}$ for $i=1,2$ (since they will be contained in the coset $g(G_{v_{1}}\cap G_{v_{2}})$, which is finite). Call this set $F$.
By Lemma˜5.17, up to taking a power of $\Phi$, the isometry $F_{\Phi}$ fixes $v_{1},gv_{1},v_{2}$ and $gv_{2}$. Thus, for $i=1,2$ we have that $\Phi$ preserves the set of elements sending $v_{i}$ to $gv_{i}$, and in particular their finite intersection $F$. As $F$ is finite and contains $g$, a power of $\Phi$ fixes $g$. This shows that $g\in H=\operatorname{Per}(\Phi)$ (recalling that the periodic subgroup is invariant under taking powers) and $N(H)=H$. ∎
Theorem 5.21.
Let $G$ be a hyperbolic group relative to $\mathcal{P}$}. Let ${\mathcal{H}}_{0}$ be a finite set of conjugacy classes of finitely generated subgroups of $G$ such that G is one-ended relative to ${\mathcal{P}}\cup{\mathcal{H}}_{0}$. Let $\phi=[\Phi]\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}}_{0}^{(t)})$. Let $\phi^{N}$ be an almost rotationless power of $\phi$.
For every rigid vertex $v\in VT^{\mathrm{Per}}$, there exists $H]\in\mathrm{Per}_{\mathrm{NP}}(\phi^{N})$ such that $G_{v}=H$. Conversely, for every $H]\in\mathrm{Per}_{\mathrm{NP}}(\phi^{N})$, there exists a rigid vertex $v\in VT^{\mathrm{Per}}$ with $G_{v}=H$.
@@
Up to taking a power of $\phi$, we may assume that $\phi$ is almost rotationless (see Remark˜5.16). Let $v\in VT^{\mathrm{Per}}$ be rigid. We claim that there exists $H]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$ with $H\leq G_{v}$. Consider the action of $G_{v}$ in $T^{\mathrm{ref}}$. If $G_{v}$ fixes a vertex in $T^{\mathrm{ref}}$, then $G_{v}$ also stabilises a non-elementary vertex $w$ of $T^{\phi}$. If $w$ is rigid, then $\langle\phi\rangle^{0}\to\operatorname{\mathrm{Out}}(G_{w})$ has finite image, so $G_{v}$ is periodic for some $\phi^{k}$, and so for $\phi$ (since we assume it is almost rotationless).
@@
Otherwise $w$ is a QH with fibre vertex. Note that all elementary subgroups of $G_{w}$ are virtually cyclic. In particular, the edges adjacent to $v\in VT^{\mathrm{Per}}$ are all virtually cyclic, hence finitely generated. Thus, $G_{v}$ is a finitely generated subgroup of $G_{w}$. Since $w$ is a QH with fibre vertex, it is locally quasi-convex. Thus, $G_{v}$ is hyperbolic.
@@
Note that, since $G_{v}$ is hyperbolic and a subgroup of $\operatorname{\mathrm{Stab}}(w)$, the restriction ${\mathcal{P}}\cup{\mathcal{H}}_{0}\cup\mathrm{Per}_{\mathrm{NP}}(\phi)|_{G_{v}}$ of ${\mathcal{P}}\cup{\mathcal{H}}_{0}\cup\mathrm{Per}_{\mathrm{NP}}(\phi)$ to $G_{v}$ is a finite family of virtually cyclic subgroups of $G_{v}$ by Lemma˜5.2. Thus, we can apply [27, Theorem 3.9] to show that $\operatorname{\mathrm{Out}}(G_{v},{\mathcal{P}}\cup{\mathcal{H}}_{0}^{(t)}\cup\mathrm{Per}_{\mathrm{NP}}(\phi)|_{G_{v}}\cup\mathrm{Inc}_{v})$ (which has $\operatorname{\mathrm{Out}}(G_{v},{\mathcal{P}}\cup({\mathcal{H}}_{0}\cup\mathrm{Per}_{\mathrm{NP}}(\phi)|_{G_{v}}\cup\mathrm{Inc}_{v})^{(t)})$ as a finite index subgroup) is infinite if and only if $G_{v}$ has an $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup\mathrm{Per}_{\mathrm{NP}}(\phi)|_{G_{v}}\cup\mathrm{Inc}_{v})$-splitting. Since $v$ is rigid, no such splitting of $G_{v}$ exists. Therefore, $\operatorname{\mathrm{Out}}(G_{v},{\mathcal{P}}\cup{\mathcal{H}}_{0}^{(t)}\cup\mathrm{Per}_{\mathrm{NP}}(\phi)|_{G_{v}}\cup\mathrm{Inc}_{v})$ is finite. In particular, the group $G_{v}$ is a nonelementary periodic subgroup of some power of $\phi$.
@@
Now suppose that $G_{v}$ is not elliptic in $T^{\mathrm{ref}}$. By Lemma˜4.3, $G_{v}$ contains a loxodromic element in $T^{\mathrm{ref}}$ and the minimal invariant tree $T^{\mathrm{ref}}_{G_{v}}\subseteq p_{\operatorname{Per}}^{-1}(v)$ of $G_{v}$ in $T^{\mathrm{ref}}$ is infinite. If this is contained in $T^{\mathrm{ref}}_{H}$ for some $H\in\mathrm{Per}_{\mathrm{NP}}(\phi)$, then $v$ is the unique vertex of $T^{\mathrm{Per}}$ stabilised by $H$, and by Lemma˜5.20 $G_{v}$ (as the stabiliser of this vertex) is equal to $H$.
@@
So suppose $T^{\mathrm{ref}}_{G_{v}}$ is not contained in any $T^{\mathrm{ref}}_{H}$. Since the $T^{\mathrm{ref}}_{H}$ are disjoint by ˜5.18, there is some edge of $T^{\mathrm{ref}}_{G_{v}}$ contained in no $T^{\mathrm{ref}}_{H}$, and hence collapsing all the $T^{\mathrm{ref}}_{H}$ will give an $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup\mathrm{Per}_{\mathrm{NP}}(\phi))$-tree where $G_{v}$ is not elliptic, contradicting ˜4.2 (7).
@@
Conversely, let $H]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$. By construction of $T^{\mathrm{Per}}$, the group $H$ fixes a vertex $v$ of $T^{\mathrm{Per}}$. Since $H$ is nonelementary, such a vertex is unique and is rigid by ˜4.2 $(10)$. By Lemma 5.20, we have $G_{v}=H$. ∎
@@
Let $G$ be a hyperbolic group relative to $\mathcal{P}$}. Let ${\mathcal{H}}_{0}$ be a finite set of conjugacy classes of finitely generated subgroups of $G$ such that G is one-ended relative to ${\mathcal{P}}\cup{\mathcal{H}}_{0}$. Let $\phi=[\Phi]\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}}_{0}^{(t)})$. Let ${\mathcal{H}}=\bigcup_{n\in\mathbb{N}^{*}}\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$.
@@
Let $\gamma$ be a geodesic edge path of length $3$ in $T^{\mathrm{Per}}$ preserved by an automorphism $\Psi\in\phi$. Suppose that there exists an edge $e$ of $\gamma$ whose stabiliser is virtually cyclic (this applies in particular when one of the vertices of $\gamma$ is QH with fibre). Hence $\operatorname{\mathrm{Out}}(G_{e})$ is finite. Thus, $\Psi$ has a power acting as the identity on an infinite cyclic subgroup of $G_{e}$.
Thus, we may suppose that $\gamma$ only contains elementary and rigid vertices. Since $\gamma$ has length $3$ and since $T^{\mathrm{Per}}$ is bipartite, it contains an interior vertex $v$ which is rigid. Let $e_{1},e_{2}$ be the two edges of $\gamma$ adjacent to $v$. Then $G_{e_{1}}\cap G_{e_{2}}$ is finite by ˜4.2 $(8)$ and for every $i\in\{1,2\}$ the group $G_{e_{i}}$ is its own normaliser in $G_{v}$.
Let $\phi^{N}$ be an almost rotationless power of $\phi$. By ˜5.21, there exists $H]\in\mathrm{Per}_{\mathrm{NP}}(\phi^{N})$ such that $G_{v}=H$. Let $\Phi_{H}\in\phi^{N}$ be such that $\mathrm{Per}(\Phi_{H})=H$. Since $\Psi$ preserves $H$, and since $H$ is its own normaliser by Lemma˜5.20, there exists $h\in H$ such that $\Psi^{N}=\mathrm{ad}_{h}\Phi_{H}$. As $\Psi^{N}$ preserves both $G_{e_{1}}$ and $G_{e_{2}}$, we see that $h\in N(G_{e_{1}})\cap N(G_{e_{2}})$, which is finite as explained above. Up to taking a power of $\Phi_{H}$ and $\Psi^{N}$, we may assume that both $\Phi_{H}$ and $\Psi^{N}$ act trivially on $N(G_{e_{1}})\cap N(G_{e_{2}})$. Taking further powers of $\Psi^{N}$ and $\Phi_{H}$ shows that there exists $M\in\mathbb{N}$ such that $\Psi^{M}=\Phi_{H}^{M}$.
Let $g\in H$ be of infinite order, which exists since $H$ is nonelementary. Then $\Psi$ has a power which fixes $g$. This proves Assertion $(1)$.
@@
By Theorem 4.2 $(2)$, the action of $G$ on $T^{\mathrm{Per}}$ is acylindrical. Thus, by Lemma 4.5 $(2)$ (which we can apply by Assertion $(1)$), it suffices to prove that, for every $n\in\mathbb{N}^{*}$, every $\Psi\in\phi^{n}$ and every $g\in\operatorname{\mathrm{Fix}}(\Psi)$, the element $g$ is elliptic in $T$.
@@
Suppose now that $g$ is nonperipheral. Suppose towards a contradiction that $g$ is loxodromic in $T^{\mathrm{Per}}$. Since $\Psi(g)=g$, the characteristic set of the isometry $F_{\Psi}$ contains the axis of $g$. By Lemma˜4.4, up to taking a power of $\phi$ and changing the representative $\Psi$, the isometry $F_{\Psi}$ fixes pointwise the axis of $g$. By Lemma 4.5 $(1)$, up to taking a power of $\Psi$, the automorphism $\Psi$ fixes elementwise a nonabelian free group of loxodromic elements. In particular, since every peripheral element fixes a point in $T^{\mathrm{Per}}$, we see that $\Psi$ fixes a nonabelian free group of nonperipheral elements, hence a nonelementary subgroup. Since $\phi$ is almost rotationless, for every $n\geq 1$, we have $\mathrm{Per}_{\mathrm{NP}}(\phi)=\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$, so that $\mathrm{Per}(\Psi)]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$. Thus, there exists $H]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$ such that $g\in\operatorname{\mathrm{Fix}}(\Psi)\subseteq H$. In that case, the element $g$ is elliptic in $T^{\mathrm{Per}}$ by construction of $T^{\mathrm{Per}}$, a contradiction.